In the displacement method to measure density, you have two objects with the same mass, the object that is less dense will:
A.) displace more water
b.) displace less water
C.) the water level will remain unchanged
D.) there is not enough information given in the problem.

I think the answer is D. But I’m not entirely sure
1
Expert's answer
2020-08-28T08:07:00-0400

A.) displace more water (Answer)

Explanation:

density(d) = mass(m)/volume(v)

d1=m/v1 and d2=m/v2

d1/d2 = v2/v1

When d1 greater than d2

Then v2 greater than v1

We know that, a submerge object displaces a volume of liquid equal to the volume of the object.

Hence, the object that is less dense has larger volume and displaces more water.